OpenStudy (anonymous):

FAN AND MEDAL! How do I do this, or what is the answer? The length of the second side of a triangle is 24 cm shorter than twice the length of the first side. Using f for the first side and s for the second side, translate the sentence into an equation.

OpenStudy (anonymous):

ok what the question is saying to 24 * double of 24cm cause look at the key answer "twice" that means double the amount and so what is 24*2?

OpenStudy (anonymous):

48

OpenStudy (anonymous):

It's also asking me how to put it down. A. s = 2f – 24 B. s = 2f + 24 C. 2s – 24 = f D. 2s + 24=1/2f

OpenStudy (anonymous):

@Bella108

OpenStudy (anonymous):

ok how you would put it down is like this 2s+24= 1/2 cause if you double it will equal 48 BUT with that its like saying to add but you are actually multiplying but the other way round , get? so ur exact answer is D :)

OpenStudy (anonymous):

Ok thanks.

OpenStudy (anonymous):

your welcome and thx for the medal